OpenStudy (anonymous):

Indicate the equation of the line through (2, -4) and having slope of 3/5.

OpenStudy (anonymous):

_x__y=

OpenStudy (anonymous):

y = mx + b is the "generic" equation of a straight line. m = slope So m = 3/5, and you have: y = (3/5)x + b Plug in (2, -4) -4 = (3/5)(2) + b -4 = 6/5 + b -26/5 = b So the equation is: y = (3/5)x - (26/5)

OpenStudy (anonymous):

but how do I make it fit in _x__y= ? @pasido